A small manufacturer of specialty welding equipment has developed a level production plan for the next four quarters, as seen below:

Supply/Demand Info Pre-Q1   Q1 Q2 Q3 Q4   
Forecast (demand) 2,300 2,760 3,680 5,980
Regular production 3,680 3,680 3,680 3,680
Subcontract production
Ending inventory
Hired employees 11
Fired employees
Total employees 21 32 32 32 32


The table below shows additional relevant information:

Capacity Information & Cost Variables   
Production rate (units/employee/quarter) 115
Subcontractor capacity (units/quarter) 480
Regular production cost/unit $74
Holding cost/unit/quarter $14
Hiring cost/employee $980
Firing cost/employee $2,150
Subcontract cost/unit $104


What is the overall total cost for this production plan? (Display your answer to the nearest whole number.)
   

What is the total regular production cost for this production plan? (Display your answer to the nearest whole number.)
   

What is the total holding cost for this production plan? (Display your answer to the nearest whole number.)
Answer #1

Formula

What is the overall total cost for this production plan? (Display your answer to the nearest whole number.)

overall total cost = Regular production cost+Holding cost+Hiring cost

=14720*74+5980*14+11*980

= 1183780

What is the total regular production cost for this production plan? (Display your answer to the nearest whole number.)

Regular production cost = 14720*74 = 1089280

What is the total holding cost for this production plan? (Display your answer to the nearest whole number.)

Total holding cost = 5980*14 = 83720